Assume that a normal distribution of data has a mean of 20 and a standard deviation of 5. Use 68 - 95 - 99.7 rule to find the percentage of values that lie above 15.

What is the percentage of values lie above 15?
Answer #1

Solution :

Given that,

= 20

= 5

Using Empirical rule,

P( - 1< X < + 1) = 68%

P( - 2< X < + 2) = 95%

P( - 3< X < + 3) = 99.7%

P(x > 15) = 1 - P(x < 15)

= 1 - 0.16 = 0.84%